# ECM: A Unified Electronic Circuit Model for Explaining the Emergence of In-Context Learning and Chain-of-Thought in Large Language Model

### v1

@@ A.4.1 ICL Explanation
To evaluate how different embedders influence model performance, we replace the sample encoder with BERT [9], RoBERTa [22], and BGE [4]. These embeddings are used for sample retrieval in ICL to minimize potential biases that may arise if a single embedder disproportionately emphasizes specific data features. For instance, an embedder might favor samples aligned with its encoding patterns, leading to distortions in the relationship between retrieved representations and model performance. By employing diverse embedders, we aim to mitigate such biases and promote more accurate theoretical modeling for ICL.
@@
This section examines the implementation of Synthetic-CoT within the ECM framework to assess its impact on reasoning performance. Synthetic-CoT autonomously generates in-context demonstrations with chain-of-thought reasoning paths. We hypothesize that these demonstrations align the system’s magnetic field, reducing logical resistor and enhancing overall circuit power output. To validate this, we generate a series of demonstrations with two categories by Synthetic-CoT, including:
  • Positive Magnetic Field Demonstrations: In-context demonstrations generated by Synthetic-CoT are positively aligned with the magnetic polarity of user queries direction.
  • Negative Magnetic Field Demonstrations: In-context demonstrations generated by Synthetic-CoT are with opposing magnetic polarity of user queries direction.
We analyze the magnetic field contributions of these categories, focusing on their differing effects. Subsequently, we integrate the computed power and measured accuracy to further investigate their operational mechanisms.
A.4.2 CoT Explanation
@@
We analyze the effect of direct answer output on reasoning performance through the ECM framework. This approach bypasses intermediary reasoning steps by appending the directive, “Let’s directly output the result without additional reasoning content,” which generates significant logical resistor within the system. To quantify this, we gradually improve the resistance value of reasoning resistor $R_{\text{CoT}}$, which meets largest Spearman correlations in the validation set, and employ the $R_{\text{CoT}}$ to the test set.
@@
We further investigate how Tool-Usage and Program-of-Thought (PoT) mitigate specific resistors in the reasoning process, focusing on their impact under the ECM framework on computational and planning resistors:
@@
  • Program-of-Thought (PoT): PoT targets planning resistor by employing structured programming logic to enable clearer and more efficient reasoning. This advantage stems from PoT’s superior ability to reduce planning resistor, facilitating finer-grained logical alignment. [^2] All implementation prompts come from Chen et al. [5]
These findings confirm that while Tool-Usage primarily addresses single-step computational resistor, PoT excels in optimizing overall planning resistor. Both methods significantly enhance performance within the ECM framework, highlighting their complementary contributions to reducing resistor and improving reasoning capabilities.
@@
This experiment evaluated the self-consistency strategy proposed by Wang et al. [40], which employs multiple parallel reasoning paths and a majority-vote mechanism for the final decision. This strategy is hypothesized to enhance performance by reducing the system’s resistance to diverse inputs, analogous to parallel resistors in a circuit, where total resistor decreases compared to series connections. By applying this analogy, we demonstrate how parallel reasoning can lower cognitive load, improving accuracy and efficiency in decision-making.
To examine the influence of inference scaling on model performance, we investigated the relationship between the number of parallel samples, $k$, and the metric like Accuracy or Pass@k, as described in the Inference Scaling Law. Specifically, Pass@k measures the accuracy of Coverage method which obtains a correct result when at least one sampled prediction is accurate. We modeled the reasoning process as a resistive circuit, where each sample represents a parallel resistor. As $k$ increases, the resistance of reasoning resistor approaches zero, reflecting a refined reasoning process and enhanced accuracy. Further, Coverage methods introduce zero verification resistance introduce much less overall resistance values, larger output power, and higher related performance. The experiment systematically varied $k$ from 1 to 100 in increments of 5, recording the accuracy and Pass@k metric under each condition [^3] We remove those generated samples are totally the same to keep the independent sampling.. We track individual reasoning path resistance, comparing these values with observed accuracy to validate the circuit analogy. The effectiveness of Self-consistency and Coverage methods can be mathematically proved based on the ECM detailed in Theoretical Proof.
A.5 Model Exploration Experiments
A.5.1 ICL Exploration
@@
This study examines the differing impacts of semantic diversity on static and dynamic ICL retrieval. We map all samples into the hidden space using RoBERTa. In static ICL, we employ a fixed set of demonstrations with the highest semantic diversity across the task, hypothesizing that this diversity enhances model performance by providing a wider range of contextual cues. In contrast, dynamic ICL selects demonstrations in real-time based on the test sample’s characteristics, prioritizing semantically similar examples that align closely with the input. To test these, we evaluate both retrieval strategies. For static ICL, we compare models exposed to high-diversity and random demonstration sets, measuring performance using accuracy. For dynamic ICL, we contrast random selection with retrieval strategies emphasizing semantic alignment. We hypothesize that semantic diversity improves static ICL by stabilizing decision-making, while in dynamic ICL, retrieving highly relevant demonstrations yields better accuracy. In addition, in order to prove that diversity is ineffective in the setting of dynamic ICL, we sort the semantic magnetic field strength and select as diverse samples as possible among the top-$k$ samples as ICL demonstrations.
@@
This experiment introduces demo-optimization, a novel method for optimizing demonstrations in ICL. The approach dynamically refines demonstrations based on feedback from the LLM’s outputs, which serve as gradients for improvement. Specifically, initial ICL demonstrations in prompting are selected, with both development dataset for optimization and optimization prompt configurations prepared. A feedback loop guides adjustments to the demonstrations, improving their clarity, relevance, and alignment with prompt requirements. Demonstrations are iteratively refined, and evaluated for contextual alignment with user requests in the development dataset. The method is compared against static and whole-prompt-optimized baselines, with performance assessed using task accuracy, output consistency, and learning efficiency.
A.5.2 CoT Exploration
CoT Optimization Proportion
@@ A.5.2 CoT Exploration
To evaluate the improvement in reasoning resistance $R_{i}$ to the target level of $\frac{1}{k}$, which depends heavily on the task’s reasoning demands, we conducted a meta-analysis of experimental results from diverse datasets. These datasets are divided into two main categories: (1) tasks requiring advanced planning and logical reasoning, such as mathematical derivations and multi-hop question answering, and (2) tasks reliant on extensive domain-specific knowledge, such as chemistry and physical question answering. As shown in Figure 7, our meta-analysis demonstrates that tasks requiring advanced planning gain the most from enhanced exploratory reasoning, significantly improving challenging components (e.g., a 2.8-fold optimization in planning capability) and overall task performance (e.g., a 3.6-fold optimization on AIME, a benchmark emphasizing long-term planning). In contrast, tasks relying heavily on domain-specific knowledge exhibit limited improvement without targeted optimization in those areas. This underscores the importance of domain-aware reasoning strategies. For example, tasks involving complex reasoning and knowledge integration, such as GPQA (chemistry, physics) and HotpotQA, show performance improvements directly proportional to advancements in the related knowledge domains (approximately 1:1 ratio). The effectiveness and the underlying reasons for such phenomena can be mathematically analyzed and proved based on the ECM detailed in Theoretical Proof.
@@
In this experiment, we examine how decomposing a self-consistency ensemble into smaller, incremental majority-voting steps reduces reasoning resistor in circuits, compared to the conventional single-step approach. The circuit operates at a stable supply voltage of 5 units, and reasoning resistance value and output power are evaluated as the number of steps increases. Each majority-voting step simplifies logical complexity and reduces overall resistance by incorporating less verification resistance in the decision-making process. Our findings demonstrate a consistent decline in reasoning resistance alongside a rise in output power, confirming that finer-grained majority-voting processes improve both efficiency and power output in circuits. The effectiveness of this method can be mathematically proved based on the ECM detailed in Theoretical Proof.
@@
Traditional self-consistency frameworks rely on a sequential verification model, where each step depends on the completion of the previous one. This structure often incurs high computational costs, as subsequent tasks cannot begin until earlier ones are resolved. To overcome this limitation, we introduce the “chain-of-verification (CoV)," a novel strategy that parallelizes verification tasks into independent components. Each component operates concurrently, validating distinct parts of the system without interdependence. This parallel approach reduces overall computational resistance by distributing the workload across multiple verifiers, thereby enabling more effective validation. As illustrated in Figure 6b, parallelization not only accelerates the verification process but also enhances model performance by mitigating bottlenecks. This improvement is particularly pronounced in large-scale systems, where sequential verification would otherwise impede efficiency. The effectiveness of CoV can also be mathematically proved based on the ECM detailed in Theoretical Proof.
A.6 Best Practice
In practice, we integrate all proposed strategies into the O1-preview model to address IOI, IMO problem-solving, and idea-generation tasks for much less resistance value of planning resistor.
@@ A.6 Best Practice
For the IOI task, we employ the following methods: First, using O1-preview, we maximize the optimization of the resistance value of its “planning resistor." Additionally, five high-intensity semantic field problems from the code_contest training dataset are retrieved as demonstrations, boosting the voltage supplied by additional ICL fields. Second, we apply the Coverage technique, sampling 100 instances and leveraging both provided and model-generated test cases in parallel, effectively reducing resistance in chain-of-thought reasoning. This process, executed in parallel, effectively reduces the resistance in the chain-of-thought reasoning. Furthermore, we incorporate the CoV concept to iteratively validate the final outputs, thereby optimizing the final performance. Additionally, the CoV framework is employed to iteratively validate outputs and optimize overall performance. To prevent data leakage from competition datasets in C++, competition code is generated in Python. Manual adjustments ensure compatibility between Python and C++ input/output handling, preserving the core logic for fair comparative testing.
@@
A similar approach is adopted for the IMO task. O1-preview is optimized for planning resistor, and five high-intensity semantic field problems from past IMO datasets are retrieved as demonstrations to enhance ICL field voltage. The Coverage technique is again employed, sampling 100 instances in parallel to reduce reasoning resistance. We introduce the fine-grained self-consistency approach within the CoV framework, enabling iterative validation of outputs and achieving notable performance gains. Additionally, we only evaluate answers and do not consider process scores.
@@
We utilize the AI Scientist framework [23] to generate over 10 ideas per domain and draft academic papers, limited to the abstract and introduction sections to address ethical concerns by avoiding experimental claims. A 16-member review panel, including 2 Superior reviewers (6+ years of experience), 8 Senior reviewers (4+ years), and 6 Junior reviewers (2+ years), ensures reliability. Each paper is reviewed by at least four reviewers, including two Senior or Superior reviewers, following NeurIPS guidelines. High-quality examples with strong semantic fields, sourced from past ICLR, NeurIPS, and CoLM openreview data, are used to guide generation. The Coverage technique, sampling of 100 instances, and iterative validation through the CoV framework further refine the process.
Figure 8: Neuron activation heatmaps for the most activated attention layer under three reasoning strategies: (a) Zero-shot CoT, (b) ICL with Semantic Magnetic Field Retrieval, and (c) Program of Thought (PoT).
A.7 Internal Mechanism underlying the ECM
To investigate the internal mechanisms underlying ECM, we conduct experiments in the LLaMA3-8B [12] focusing on three strategies: Zero-shot CoT, ICL with Semantic Magnetic Field Retrieval, and PoT. By analyzing neuron activation values and frequencies—particularly in the most activated attention layer—we aim to elucidate the relationship between reasoning strategies and neural dynamics. These findings support the theoretical framework of ECM. The experiments use the BigGSM dataset, a standard benchmark for mathematical reasoning tasks. For ICL, demonstrations are dynamically retrieved using the BGE model in a 3-shot configuration. Figure 8 shows neuron activation heatmaps for the most activated attention layer under each reasoning strategy. For Zero-shot CoT, the heatmap reveals sparse activations, with limited regions of intense activity, highlighting the difficulty of reasoning without contextual support. Conversely, ICL and PoT display more distributed and elevated activation patterns. These suggest that the underlying reasoning mechanisms engage a broader spectrum of neurons, with comparable activation dynamics reflecting similar reasoning power across these strategies.
@@ B.1.1 Analysis of Effectiveness of Self-consistency
First, according to the series resistor assumption of CoT, the total resistance of series resistors can be expressed as:
$R_{\text{all}}=R_{0}+R_{\text{CoT}}.$ (12)
Building upon the illustration in Figure 4c, we hypothesize that self-consistency, achieved through multiple parallel reasoning processes and subsequent result aggregation by majority voting, can be represented as an aggregation resistor integrated alongside several parallel CoT resistors. In this framework, the effective resistance of the reasoning process transitions from $R_{\text{CoT}}$ to $\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}}$, accompanied with a majority-voting resistor $R_{S}$ to determine the final outcome. Consequently, the total resistance associated with self-consistency, denoted as $R^{{}^{\prime}}_{\text{all}}$, is expressed as:
$R^{{}^{\prime}}_{\text{all}}=R_{0}+R_{S}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}},$ (13)
where $n$ is the number of samples, $R_{\text{CoT}}^{i}$ denotes the resistance corresponding to the $i-$th sample and $R_{S}$ accounts for the resistor due to self-consistency integration. This framework allows for rapid computation of the output power:
$P^{\text{self}}_{\text{out}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(R_{0}+R_{S}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}})^{2}}.$ (14)
As the number of samples increases, the curves in Figure 4c reveal that the real performance curve is almost exactly consistent with the theoretical range. Furthermore, assume that $\{R_{\text{CoT}}^{i}\}_{i=1}^{\infty}$ is bounded thus we can easily prove that the overall resistance has a lower bound:
$\lim_{n\rightarrow+\infty}R^{\prime}_{\text{all}}=R_{0}+R_{S}+\lim_{n\rightarrow+\infty}\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{CoT}}^{i}}}=R_{0}+R_{S}.$ (15)
@@
$\hat{P}^{\text{self}}_{\text{out}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}\right)^{2}R_{0}}{(R_{0}+R_{S})^{2}}$ (16)
As shown in Figure 4c, model performance aligns with this theoretical upper bound as sample size increases.

### v2

@@ A.4 Optimization Interpreting Experiments through Magnetic Field Adjustment
To evaluate how different encoders influence model performance, we replace the sample encoder with BERT [9], RoBERTa [27], and BGE [3]. These embeddings are used for sample retrieval in ITL to minimize potential biases that may arise if a single encoder disproportionately emphasizes specific data features. For instance, an encoder might favor samples aligned with its encoding patterns, leading to distortions in the relationship between retrieved representations and model performance. By employing diverse encoders, we aim to mitigate such biases and promote more accurate theoretical modeling for ITL.
Meta-Recognition Representations
In this experiments, all demonstration representation methods employ the meta-recognition [10] framework by default. Specifically, GPT-4o [1] is utilized to extract sub-capability tags, and then RoBERTa [27] concatenates and represents these tags in the hidden space.
@@
This section examines the implementation of Synthetic-CoT within the ECP framework to assess its impact on reasoning performance. Synthetic-CoT autonomously generates in-context demonstrations with Inference-Time Reasoning reasoning paths. We hypothesize that these demonstrations align the system’s magnetic field, reducing logical resistance and enhancing overall circuit power output. To validate this, we generate a series of demonstrations by Synthetic-CoT. Specifically, In-context demonstrations generated by Synthetic-CoT are positively aligned with the magnetic polarity of user queries direction. We analyze the magnetic field contributions of these categories, focusing on their differing effects. Subsequently, we integrate the computed power and measured accuracy to further investigate their operational mechanisms.
A.5 Optimization Interpreting Experiments through Resistor Reorganization
@@
We analyze the effect of direct answer output on reasoning performance through the ECP framework. This approach bypasses intermediary reasoning steps by appending the directive, “Let’s directly output the result without additional reasoning content,” which generates a significant logical resistor within the system. To quantify this, we gradually improve the resistance value of reasoning resistor $R_{\text{ITR}}$, which meets the largest Spearman correlations in the validation set, and employ the $R_{\text{ITR}}$ to the test set.
@@
We further investigate how Tool-Usage and Program-of-Thought (PoT) mitigate specific resistors in the reasoning process, focusing on their impact under the ECP framework on computational and planning resistors:
@@
  • Program-of-Thought (PoT): PoT targets planning resistor by employing structured programming logic to enable clearer and more efficient reasoning. This advantage stems from PoT’s superior ability to reduce planning resistor, facilitating finer-grained logical alignment. [^2] All implementation prompts come from Chen et al. [4].
These findings confirm that while Tool-Usage primarily addresses the single-step computational resistor, PoT further excels in optimizing the overall planning resistor. Both methods significantly enhance performance within the ECP framework, highlighting their complementary contributions to reducing resistance and improving reasoning capabilities.
@@
This experiment evaluated the self-consistency strategy proposed by Wang et al. [41], which employs multiple parallel reasoning paths and a majority-vote mechanism for the final decision. This strategy is hypothesized to enhance performance by reducing the system’s resistance to diverse inputs, analogous to parallel resistors in a circuit, where the total resistance decreases compared to series connections. By applying this analogy, we demonstrate how parallel reasoning can lower cognitive load, improving accuracy and efficiency in decision-making.
To examine the influence of inference scaling on model performance, we investigated the relationship between the number of parallel samples, $k$, and the metric like Accuracy or Pass@k, as described in the Inference Scaling Law. Specifically, Pass@k measures the accuracy of the Coverage method which obtains a correct result when at least one sampled prediction is accurate. We modeled the reasoning process as a resistive circuit, where each sample represents a parallel resistor. As $k$ increases, the resistance of reasoning resistor approaches zero, reflecting a refined reasoning process and enhanced accuracy. Further, Coverage methods introduce zero verification resistance introduce much less overall resistance values, larger output power, and higher related performance. The experiment systematically varied $k$ from 1 to 100 in increments of 5, recording the accuracy and Pass@k metric under each condition. We track individual reasoning path resistance, comparing these values with observed accuracy to validate the circuit analogy. The effectiveness of Self-consistency and Coverage methods can be mathematically proved based on the ECP detailed in Theoretical Proof.
We annotated 100 instances, each comprising 32 outputs sampled from the model, with four annotators involved in the labeling. For each instance, the 32 outputs were presented to the annotators in sequence; they were asked to indicate whether each output contained text or logical content that was exactly duplicated from any previously shown example. Finally, for sample sizes of $2^{n}$ , we counted the number of outputs exhibiting no duplication and treated these counts as independent samples.
A.6 Optimization Exploration Experiments through Magnetic Field Adjustment
@@
This study examines the differing impacts of semantic diversity on static and dynamic ITL retrieval. We map all samples into the hidden space using RoBERTa. In static ITL, we employ a fixed set of demonstrations with the highest semantic diversity across the task, hypothesizing that this diversity enhances model performance by providing a wider range of contextual cues. In contrast, dynamic ITL selects demonstrations in real-time based on the test sample’s characteristics, prioritizing semantically similar examples that align closely with the input. To test these, we evaluate both retrieval strategies. For static ITL, we compare ECP exposed to high-diversity and random demonstration sets, measuring performance using accuracy. For dynamic ITL, we contrast random selection with retrieval strategies emphasizing semantic alignment. We hypothesize that semantic diversity improves static ITL by stabilizing decision-making, while in dynamic ITL, retrieving highly relevant demonstrations yields better accuracy. In addition, in order to prove that diversity is ineffective in the setting of dynamic ITL, we sort the semantic magnetic field strength and select as diverse samples as possible among the top-$k$ samples as ITL demonstrations.
@@
This experiment introduces demo-optimization, a novel method for optimizing demonstrations in ITL. The approach dynamically refines demonstrations based on feedback from the LLM’s outputs, which serve as gradients for improvement. Specifically, initial ITL demonstrations in prompting are selected, with both development dataset for optimization and optimization prompt configurations prepared. A feedback loop guides adjustments to the demonstrations, improving their clarity, relevance, and alignment with prompt requirements. Demonstrations are iteratively refined, and evaluated for contextual alignment with user requests in the development dataset. The method is compared against static and whole-prompt-optimized baselines, with performance assessed using task accuracy, output consistency, and learning efficiency.
A.7 Optimization Exploration Experiments through Resistor Reorganization
ITR Optimization Proportion
@@ A.7 Optimization Exploration Experiments through Resistor Reorganization
To evaluate the improvement in reasoning resistance $R_{i}$ to the target level of $\frac{1}{k}$, which depends heavily on the task’s reasoning demands, we conducted a meta-analysis of experimental results from diverse datasets. These datasets are divided into two main categories: (1) tasks requiring advanced planning and logical reasoning, such as mathematical derivations and multi-hop question answering, and (2) tasks reliant on extensive domain-specific knowledge, such as chemistry and physical question answering. As shown in Fig. 10, our meta-analysis demonstrates that tasks requiring advanced planning gain the most from enhanced exploratory reasoning, significantly improving challenging components (e.g., a 2.8-fold optimization in planning capability) and overall task performance (e.g., a 3.6-fold optimization on AIME, a benchmark emphasizing long-term planning). In contrast, tasks relying heavily on domain-specific knowledge exhibit limited improvement without targeted optimization in those areas. This underscores the importance of domain-aware reasoning strategies. For example, tasks involving complex reasoning and knowledge integration, such as GPQA (chemistry, physics) and HotpotQA, show performance improvements directly proportional to advancements in the related knowledge domains (approximately 1:1 ratio). The effectiveness and the underlying reasons for such phenomena can be mathematically analyzed and proved based on the ECP detailed in Theoretical Proof.
@@
In this experiment, we examine how decomposing a self-consistency ensemble into smaller, incremental majority-voting steps reduces the reasoning resistance in circuits, compared to the conventional single-step approach. The circuit operates at a stable supply voltage of 5 units, and reasoning resistance value and output power are evaluated as the number of steps increases. Each majority-voting step simplifies logical complexity and reduces overall resistance by incorporating less verification resistance in the decision-making process. Our findings demonstrate a consistent decline in reasoning resistance alongside a rise in output power, confirming that finer-grained majority-voting processes improve both efficiency and power output in circuits. The effectiveness of this method can be mathematically proved based on the ECP in Theoretical Proof.
@@
Traditional self-consistency frameworks rely on a sequential verification, where each step depends on the completion of the previous one. This structure often incurs high computational costs, as subsequent tasks cannot begin until earlier ones are resolved. To overcome this limitation, we introduce the “chain-of-verification (CoV)," a novel strategy that parallelizes verification tasks into independent components. Each component operates concurrently, validating distinct parts of the system without interdependence. This parallel approach reduces overall computational resistance by distributing the workload across multiple verifiers, thereby enabling more effective validation. As illustrated in Fig. 7b, parallelization not only accelerates the verification process but also enhances model performance by mitigating bottlenecks. This improvement is particularly pronounced in large-scale systems, where sequential verification would otherwise impede efficiency. The effectiveness of CoV can also be mathematically proved in Theoretical Proof.
A.8 Best Practice
In practice, we integrate all proposed strategies into the o1-preview model to address IOI, IMO, and idea-generation tasks for much less resistance value of planning resistor.
@@ A.8 Best Practice
For the IOI task, we employ the following methods: First, using o1-preview, we maximize the optimization of the resistance value of its “planning resistor." Additionally, five high-intensity semantic field problems from the code_contest training dataset are retrieved as demonstrations, boosting the voltage supplied by additional ITL fields. Second, we apply the Coverage technique, sampling 100 instances and leveraging both provided and model-generated test cases in parallel, effectively reducing resistance in Inference-Time Reasoning reasoning. This process, executed in parallel, effectively reduces the resistance in the Inference-Time Reasoning reasoning. Furthermore, we incorporate the CoV concept to iteratively validate the final outputs, thereby optimizing the final performance. Additionally, the CoV framework is employed to iteratively validate outputs and optimize overall performance. To prevent data leakage from competition datasets in C++, competition code is generated in Python. Manual adjustments ensure compatibility between Python and C++ input/output handling, preserving the core logic for fair comparative testing.
@@
A similar approach is adopted for the IMO task. o1-preview is optimized for planning resistor, and five high-intensity semantic field problems from past IMO datasets are retrieved as demonstrations to enhance ITL field voltage. The Coverage technique is again employed, sampling 100 instances in parallel to reduce reasoning resistance. We introduce the fine-grained self-consistency approach within the CoV framework, enabling iterative validation of outputs and achieving notable performance gains. Additionally, we only evaluate answers and do not consider process scores.
@@
We utilize the AI Scientist framework [28] to generate over 10 ideas per domain and draft academic papers, limited to the abstract and introduction sections to address ethical concerns by avoiding experimental claims. A 16-member review panel, including 2 Superior reviewers (6+ years of experience), 8 Senior reviewers (4+ years), and 6 Junior reviewers (2+ years), ensures reliability. Each paper is reviewed by at least four reviewers, including two Senior or Superior reviewers, following NeurIPS guidelines. High-quality examples with strong semantic fields, sourced from past ITLR, NeurIPS, and CoLM openreview data, are used to guide generation. The Coverage technique, sampling of 100 instances, and iterative validation through the CoV framework further refine the process.
Figure 11: Neuron activation heatmaps for the most activated attention layer under three reasoning strategies: (a) Zero-shot ITR, (b) ITL with Semantic Magnetic Field Retrieval, and (c) Program of Thought (PoT).
A.9 Internal Mechanism underlying the ECP
To investigate the internal mechanisms underlying ECP, we conduct experiments in the LLaMA3-8B [13] focusing on three strategies: Zero-shot ITR, ITL with Semantic Magnetic Field Retrieval, and PoT. By analyzing neuron activation values and frequencies, particularly in the most activated attention layer, we aim to elucidate the relationship between reasoning strategies and neural dynamics. These findings support the theoretical framework of ECP. The experiments use the BigGSM dataset, a standard benchmark for mathematical reasoning tasks. For ITL, demonstrations are dynamically retrieved using the BGE model in a 3-shot configuration. Fig. 11 shows neuron activation heatmaps for the most activated attention layer under each reasoning strategy. For Zero-shot ITR, the heatmap reveals sparse activations, with limited regions of intense activity, highlighting the difficulty of reasoning without contextual support. Conversely, ITL and PoT display more distributed and elevated activation patterns. These suggest that the underlying reasoning mechanisms engage a broader spectrum of neurons, with comparable activation dynamics reflecting similar reasoning power across these strategies.
@@ B.1.1 Analysis of Effectiveness of Self-consistency
First, according to the series resistor assumption of ITR, the total resistance of series resistors can be expressed as:
$R_{\text{all}}=R_{0}+R_{\text{ITR}}.$ (14)
Building upon the illustration in Fig. 5c, we hypothesize that self-consistency, achieved through multiple parallel reasoning processes and subsequent result aggregation by majority voting, can be represented as an aggregation resistor integrated alongside several parallel ITR resistors. In this framework, the effective resistance of the reasoning process transitions from $R_{\text{ITR}}$ to $\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{ITR}}^{i}}}$, accompanied with a majority-voting resistor $R_{S}$ to determine the final outcome. Consequently, the total resistance associated with self-consistency, denoted as $R^{{}^{\prime}}_{\text{all}}$, is expressed as:
$R^{{}^{\prime}}_{\text{all}}=R_{0}+R_{S}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{ITR}}^{i}}},$ (15)
where $n$ is the number of samples, $R_{\text{ITR}}^{i}$ denotes the resistance corresponding to the $i$-th sample and $R_{S}$ accounts for the resistor due to self-consistency integration. This framework allows for rapid computation of the output power:
$P^{\text{self}}_{\text{out}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}R_{0}}{(R_{0}+R_{S}+\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{ITR}}^{i}}})^{2}}.$ (16)
As the number of samples increases, the curves in Fig. 5c reveal that the real performance curve is almost exactly consistent with the theoretical range. Furthermore, assume that $\{R_{\text{ITR}}^{i}\}_{i=1}^{\infty}$ is bounded thus we can easily prove that the overall resistance has a lower bound:
$\lim_{n\rightarrow+\infty}R^{\prime}_{\text{all}}=R_{0}+R_{S}+\lim_{n\rightarrow+\infty}\frac{1}{\sum_{i=1}^{n}\frac{1}{R_{\text{ITR}}^{i}}}=R_{0}+R_{S}.$ (17)
@@
$\hat{P}^{\text{self}}_{\text{out}}=\frac{\left(\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}\right)^{2}R_{0}}{(R_{0}+R_{S})^{2}}$ (18)
As shown in Fig. 5c, model performance aligns with this theoretical upper bound as sample size increases.